# appendices

### v1

@@ Abstract
Textual descriptions in cyber threat intelligence (CTI) reports, such as security articles and news, are rich sources of knowledge about cyber threats, crucial for organizations to stay informed about the rapidly evolving threat landscape. However, current CTI extraction methods lack flexibility and generalizability, often resulting in inaccurate and incomplete knowledge extraction. Syntax parsing relies on fixed rules and dictionaries, while model fine-tuning requires large annotated datasets, making both paradigms challenging to adapt to new threats and ontologies. To bridge the gap, we propose CTINexus, a novel framework leveraging optimized in-context learning (ICL) of large language models (LLMs) for data-efficient CTI knowledge extraction and high-quality cybersecurity knowledge graph (CSKG) construction. Unlike existing methods, CTINexus requires neither extensive data nor parameter tuning and can adapt to various ontologies with minimal annotated examples. This is achieved through: (1) a carefully designed automatic prompt construction strategy with optimal demonstration retrieval for extracting a wide range of cybersecurity entities and relations; (2) a hierarchical entity alignment technique that canonicalizes the extracted knowledge and removes redundancy; (3) an ICL-enhanced long-distance relation prediction technique to further complete the CKSG with missing links. Our extensive evaluations using 150 real-world CTI reports collected from 10 platforms demonstrate that CTINexus significantly outperforms existing methods in constructing accurate and complete CSKGs, highlighting its potential to transform CTI analysis with an efficient and adaptable solution for the dynamic threat landscape.
@@
Threat Intelligence, Large Language Model, In-Context Learning, Cybersecurity Knowledge Graph
@@ 1 Introduction
Modern cyberattacks are becoming increasingly complex and rapidly evolving. Many public and commercial organizations extensively record and share cyber threat intelligence (CTI) on their platforms to combat evolving threats. According to Gartner, CTI is defined as “evidence-based knowledge, including context, mechanisms, indicators, implications and actionable advice, about an existing or emerging menace or hazard to assets that can be used to inform decisions regarding the subject’s response to that menace or hazard” [60]. Such knowledge is crucial for organizations to monitor the rapidly evolving threat landscape, promptly detect early signs of an attack, and effectively contain the attack with proper measures. Given its importance, CTI has been increasingly collected and exchanged across organizations, often in the form of Indicators of Compromise (IOC) [56]. IOCs are forensic artifacts of an intrusion such as virus signatures, IPs/domains of botnets, MD5 hashes of attack files, etc. However, recent studies [56, 77] showed that knowledge offered by IOCs is rather limited, which covers only a limited set of knowledge and has a short lifespan.
@@
Recognizing the limitations of IOCs, recent research has shifted towards automatically extracting richer knowledge from textual threat descriptions in CTI reports (i.e., CTI text). These reports, such as security blog articles [22, 6] and news [10, 20], are produced by security researchers and practitioners and published on websites, summarizing threat behaviors in natural language. Besides IOCs, these reports contain various other cybersecurity entities, such as malware, vulnerabilities, and attack techniques, as well as their relationships, illustrating their interactions and dependencies. This knowledge is crucial for building a comprehensive cyber threat profile.
@@
Several approaches have been proposed for automatically extracting security knowledge from CTI text and constructing a cybersecurity knowledge graph (CSKG). Syntax parsing-based approaches [43, 39, 56] leverage fixed dependency rules and hand-crafted dictionaries to parse the grammatical structure of sentences and extract key subject-verb-object triplets. Fine-tuning-based approaches [71, 55, 30] leverage pre-trained transformer models and fine-tune them on labeled CTI text datasets to identify semantic roles and extract entities and relations. However, all these existing methods suffer from several key drawbacks (see Section 2.2 for details), particularly when facing the evolving threat landscape: (1) Lack of flexibility and generalizability: Many of these methods are tailored to specific cybersecurity ontologies, focusing on a fixed set of entities and relation types. They are difficult to generalize to new ontologies and emerging threats and terminologies. Fixed rules have limited flexibility to adapt to new patterns and require manual creation and maintenance. Model fine-tuning, however, requires a large amount of labeled CTI data. Such data is scarce in security, especially for new threats that lack annotations. (2) Information inaccuracy and incompleteness: Due to the peculiarities of the security context and the lack of deep analysis, these methods often generate low-quality CSKGs that are incomplete, inaccurate, and disconnected. Fig. 1 shows example CSKGs generated by two representative methods for a real-world CTI report. We can observe several issues, including incomplete entities, misidentified entity boundaries, misaligned entities, missing links, etc. These low-quality CSKGs limit the ability to obtain a comprehensive threat profile, hindering the effective use of CTI to enhance other defensive measures.
@@
We present CTINexus, an LLM-powered framework for automated CTI knowledge extraction and CSKG construction from CTI reports. Unlike existing methods limited by generalizability and data demands, CTINexus introduces an optimized-ICL-based pipeline for data-efficient inference, enabling precise extraction of diverse cybersecurity entities and relations while adapting to various ontologies. By leveraging an optimal ICL strategy, CTINexus generalizes from selected demonstrations to perform versatile CSKG construction tasks. In addition to cybersecurity triplet extraction, CTINexus also refines extracted knowledge to enhance the canonicalization and completeness of the resulting knowledge graph. As shown in Fig. 1, the CSKG constructed by CTINexus has significantly higher quality compared to existing approaches.
@@
To further complete the CSKG with implicit relations for distant entities, we designed a long-distance relation prediction technique. First, entities with the highest degree centrality in a subgraph are selected as the central nodes of that subgraph. Then, CTINexus performs ICL-enhanced implicit relation prediction on these central nodes, conditioned on the input context, to infer connections among the disjoint subgraphs.
@@
We conducted comprehensive evaluations using 150 CTI reports from 10 well-recognized CTI sharing platforms [3, 4, 6, 9, 20, 21, 10, 22, 23, 25]. CTINexus achieved F1 scores of 87.65% in cybersecurity triplet extraction, 89.94% in coarse-grained entity grouping, 99.80% in fine-grained entity merging, and 90.99% in long-distance relation prediction. Qualitative analysis showed that CTINexus constructs more comprehensive and interconnected CSKGs compared to TTPDrill [43], EXTRACTOR [71], and LADDER [30]. Quantitatively, CTINexus outperforms EXTRACTOR by 25.36% in F1 score for cybersecurity triplet extraction and LADDER by 19% in cybersecurity entity extraction. We also explored various prompting strategies and four backbone models (closed-source models: GPT-3.5, GPT-4; open-source models: Llama3 and QWen2.5) to identify the optimal ICL paradigm for CTI knowledge extraction, providing valuable insights for future research.
@@ 2.1 Cyber Threat Intelligence
Although crowd-sourced CTI reports provide valuable information, their unstructured format significantly hinders their effectiveness. As the number and complexity of cyberattacks increase, the textual CTI descriptions have also expanded, creating an urgent need for automated information extraction from CTI [68]. The extracted knowledge can be used to construct cybersecurity knowledge graphs (CSKGs), where nodes represent entities and edges represent relations. Compared to unstructured CTI text, CSKGs provide a holistic profile for cyber threats, offer better visualization, and are more amenable to integration into downstream applications. The construction of a CSKG typically follows an ontology, which specifies the entity types and their allowed relations. However, despite efforts to create multiple security ontologies [44, 69, 75] covering different aspects of threats, keeping up with the rapidly evolving threat landscape remains challenging as new threats, techniques, and tools constantly emerge. It is nearly impossible to develop a universal ontology that encompasses all current and future threats. This underscores the need for CTI knowledge extraction approaches that can flexibly adapt to different ontologies and emerging threats with minimal transition effort.
@@ 2.2 Limitations of Existing Approaches
  • Domain complexity: The grammatical rules can apply to any domain. However, CTI text in the security domain has several peculiarities that can confuse syntax parsing, leading to inaccurate extraction. Cybersecurity entities can contain special characters, such as dots in IPv4 addresses, underscores in file names, and slashes in file paths. These special characters can confuse basic NLP modules, like sentence segmentation and tokenization, which syntax parsing relies upon.
@@
On the other hand, fine-tuning-based approaches fine-tune pre-trained neural networks on annotated CTI domain datasets to perform named entity recognition (NER) and relation extraction (RE). For example, EXTRACTOR [71] fine-tunes a pre-trained BERT [72] model with thousands of annotated CTI sentences, to perform semantic role labeling to extract subjects, objects, and verb actions. AttacKG [55] fine-tunes a pre-trained model in the SpaCy library [41] to recognize entities and extract dependencies. LADDER [30] fine-tunes different pre-trained transformer models, including BERT, RoBERTa, and XML-RoBERTa, on their custom datasets annotated according to their own ontology for performing NER and RE. However, fine-tuning-based methods also have several drawbacks:
  • Resource requirement: Fine-tuning requires large amounts of labeled data (i.e., annotated CTI text corpora), and the labeling needs to be aligned with the targeted ontology. Such annotations are expensive to obtain, especially for emerging threats. Additionally, fine-tuning can be computationally expensive if the backbone model contains lots of parameters.
  • Ontology lock-in: Since the models are fine-tuned on datasets annotated using a specific ontology, they are difficult to generalize to new ontologies that cover different entities and relations. Transferring to other ontologies would require reannotating vast data and retraining the models.
@@ 2.2.1 Motivating Example
As shown in Fig. 1, the CSKG constructed by CTINexus is comprehensive, well-connected, and of much higher quality, addressing all previous drawbacks. By leveraging the in-context learning of LLMs, the construction of such a CSKG does not rely on large amounts of training data and can flexibly adapt to different ontologies. We describe our approach in Section 4.
Figure 2: Overview of CTINexus. CTINexus comprises three phases. Phase 1, Security Triplet Extraction, enables end-to-end extraction of cybersecurity triplets using in-context learning of LLM. Phase 2, Hierarchical Entity Alignment, reduces the redundancy of CSKG through coarse-grained grouping and fine-grained clustering. Phase 3, Long-Distance Relation Prediction, connects disjoint subgraphs by identifying central nodes and performing relation inference.
2.3 Large Language Models
@@ 3 Overview
Fig. 2 illustrates CTINexus. CTINexus introduces a novel ICL-based approach for data-efficient CTI knowledge extraction and CSKG construction. Unlike previous methods, CTINexus eliminates the need for extensive data annotations and parameter tuning, facilitating easy generalization to different ontologies. CTINexus focuses on constructing a connected and comprehensive CSKG, enabling entity alignment and long-distance relation inference. CTINexus consists of three phases.
@@ 4.2 Cybersecurity Triplet Extraction
Figure 3: Comparison of CTINexus’s ICL-based CTI knowledge extraction (left) and a multi-turn QA-based extraction (right). CTINexus consolidates task descriptions (including applied ontology), $k$ selected demonstrations, and query into a single instruction for efficient cybersecurity triplet extraction. In contrast, the multi-turn QA paradigm requires multiple rounds of conversations with multiple prompts to extract different entities and relations, resulting in inefficiency.
Figure 4: The design of CTINexus’s hierarchical entity alignment. The coarse-grained entity grouping module populates an ICL prompt to assign entity types to the extracted triplets according to the applied ontology. Entities with the same type are grouped together. The fine-grained entity merging module then uses an embedding-based technique to merge semantically similar entities within each group based on a predefined similarity threshold.
Given that CTI text may contain diverse relations and we want the approach to be adaptable to emerging threats, we formulate the cybersecurity triplet extraction module in our pipeline as a semi-open extraction problem: Entity types are prescribed using MALOnt, as its coverage is already comprehensive, while relation extraction is modeled as open RE to maximize the coverage. These approaches transform information extraction tasks into multi-turn question-answering, leveraging the conversational capabilities of LLMs. Fig. 3 illustrates this paradigm. This method involves creating multiple questioning prompts for each information type and refining the responses. However, applying this multi-turn QA formulation to cybersecurity entity and relation extraction requires numerous lengthy prompts due to the extensive cybersecurity ontology that could contain many entity classes. For $N$ entities in the input CTI, $\frac{N(N-1)}{2}$ prompts are needed to extract relations between identified entities, leading to repetitive content and significant token waste, hindering scalability. Additionally, the multi-turn paradigm suffers from confirmation bias [34], as LLMs may confirm with a non-existing relation after several rounds of dialogue. These erroneous links can be particularly harmful in the CTI domain, negatively affecting downstream defense solutions by producing false alarms.
@@
    Demonstrations: Top-$k$ most relevant examples are retrieved using the demonstration retriever. Each example consists of a CTI report annotated with the security triplets. These examples are ordered in ascending similarity to the input query based on findings described in Section 5.3.
@@ 4.3 Hierarchical Entity Alignment
Entity alignment identifies entities with different mentions that refer to the same real-world object, a key area in knowledge graph research [26]. Aligning these mentions integrates sub-graphs containing complementary knowledge, enhancing the comprehensiveness of the knowledge graph. Traditional entity alignment techniques rely on heuristics like string matching and structural similarities, which fail to capture the underlying semantics or context of entities and have limited accuracy.
Recent studies [74, 78, 65] have adopted deep learning-based methods to learn vector representations (i.e., embeddings) of entities, achieving better accuracy. However, embedding-based techniques face unique challenges in our problem domain. In CTI text, entities with similar embeddings may refer to different concepts, e.g., “.akira files” (an IOC) and “Akira” (a threat actor). Besides, comparing the semantic distance between every pair of entities has a computational complexity of $n^{2}$, where $n$ is the total number of entities. This is inefficient when $n$ becomes large.

### v2

@@ Abstract
Textual descriptions in cyber threat intelligence (CTI) reports, such as security articles and news, are rich sources of knowledge about cyber threats, crucial for organizations to stay informed about the rapidly evolving threat landscape. However, current CTI knowledge extraction methods lack flexibility and generalizability, often resulting in inaccurate and incomplete knowledge extraction. Syntax parsing relies on fixed rules and dictionaries, while model fine-tuning requires large annotated datasets, making both paradigms challenging to adapt to new threats and ontologies. To bridge the gap, we propose CTINexus, a novel framework leveraging optimized in-context learning (ICL) of large language models (LLMs) for data-efficient CTI knowledge extraction and high-quality cybersecurity knowledge graph (CSKG) construction. Unlike existing methods, CTINexus requires neither extensive data nor parameter tuning and can adapt to various ontologies with minimal annotated examples. This is achieved through: (1) a carefully designed automatic prompt construction strategy with optimal demonstration retrieval for extracting a wide range of cybersecurity entities and relations; (2) a hierarchical entity alignment technique that canonicalizes the extracted knowledge and removes redundancy; (3) an long-distance relation prediction technique to further complete the CSKG with missing links. Our extensive evaluations using 150 real-world CTI reports collected from 10 platforms demonstrate that CTINexus significantly outperforms existing methods in constructing accurate and complete CSKG, highlighting its potential to transform CTI analysis with an efficient and adaptable solution for the dynamic threat landscape.
@@
Cyber Threat Intelligence, Large Language Model, In-Context Learning, Cybersecurity Knowledge Graph
@@ 1 Introduction
Modern cyberattacks are becoming increasingly complex and rapidly evolving. Many public and commercial organizations extensively record and share cyber threat intelligence (CTI) on their platforms to combat evolving threats. According to Gartner, CTI is defined as “evidence-based knowledge, including context, mechanisms, indicators, implications and actionable advice, about an existing or emerging threat to assets, used to inform decisions regarding the subject’s response to that threat” [7]. Such knowledge is crucial for organizations to monitor the rapidly evolving threat landscape, promptly detect early signs of an attack, and effectively contain the attack with proper measures. Given its importance, CTI has been increasingly collected and exchanged across organizations, often in the form of Indicators of Compromise (IOC) [57]. IOCs are forensic artifacts of an intrusion such as virus signatures, IPs/domains of botnets, MD5 hashes of attack files, etc. However, recent studies [57, 80] showed that knowledge offered by IOCs is rather limited, which covers only a limited set of knowledge and has a short lifespan.
@@
Recognizing the limitations of IOCs, recent research has shifted towards automatically extracting richer knowledge from textual threat descriptions in CTI reports (i.e., CTI text). These reports, such as security blog articles [25, 6] and news [11, 21], are produced by security researchers and practitioners and published on websites, summarizing threat behaviors in natural language. Besides IOCs, these reports contain various other cybersecurity entities, such as malware, vulnerabilities, and attack techniques, as well as their interactions and dependencies. This knowledge is crucial for building a comprehensive cyber threat profile.
@@
Several approaches have been proposed for automatically extracting security knowledge from CTI and constructing a cybersecurity knowledge graph (CSKG). Syntax-parsing-based approaches [48, 43, 57] leverage fixed dependency rules and hand-crafted dictionaries to parse the grammatical structure of sentences and extract key subject-verb-object triplets. Fine-tuning-based approaches [74, 56, 32] leverage pre-trained transformers and fine-tune them on labeled CTI datasets to identify semantic roles and extract entities and relations. However, these methods suffer from several key drawbacks, particularly facing the evolving threat landscape:
(1) Lack of flexibility and generalizability: Many of these methods are tailored to specific cybersecurity ontologies, focusing on a fixed set of entities and relation types. They are difficult to generalize to new ontologies and emerging threats and terminologies. Fixed rules have limited flexibility to adapt to new patterns and require manual creation and maintenance. Model fine-tuning, however, requires a large amount of labeled CTI corpus data. Such data is scarce in security, especially for new threats that lack annotations.
(2) Information inaccuracy and incompleteness: Due to the peculiarities of the security context and the lack of deep analysis, these methods often generate low-quality CSKGs that are incomplete, inaccurate, and disconnected. Fig. 1 shows example CSKGs generated by three representative methods for a real-world CTI report. We can observe issues including incomplete entities, misidentified entity boundaries, misaligned entities, missing links, etc. These low-quality CSKGs limit the ability to obtain a comprehensive threat profile, hindering the effective use of CTI to enhance defensive measures.
@@
We present CTINexus, an LLM-powered framework for automated CTI knowledge extraction and CSKG construction from CTI reports. Unlike existing methods limited by generalizability and data demands, CTINexus introduces an optimized-ICL-based pipeline for data-efficient inference, enabling precise extraction of diverse cybersecurity entities and relations while adapting to various ontologies. In addition, CTINexus refines the extracted knowledge to enhance the canonicalization and completeness of the resulting knowledge graph. As shown in Fig. 1, the CSKG constructed by CTINexus has significantly higher quality compared to existing approaches.
@@
To further complete the CSKG with implicit relations for distant entities, we designed a long-distance relation prediction technique. First, entities with the highest degree centrality in a subgraph are selected as the central nodes of that subgraph. Then, CTINexus leverages ICL to predict implicit relations among these central nodes to infer connections among the disjoint subgraphs.
@@
We conducted comprehensive evaluations using 150 CTI reports from 10 well-recognized CTI sharing platforms [3, 4, 6, 10, 21, 24, 11, 25, 26, 28]. CTINexus achieved F1-scores of 87.65% in cybersecurity triplet extraction, 89.94% in coarse-grained entity grouping, 99.80% in fine-grained entity merging, and 90.99% in relation prediction. Qualitative analysis showed that CTINexus constructs more comprehensive and interconnected CSKGs compared to TTPDrill [48], EXTRACTOR [74], and LADDER [32]. Quantitatively, CTINexus outperforms EXTRACTOR by 25.36% in F1-score for cybersecurity triplet extraction and LADDER by 19% in cybersecurity entity extraction. We also explored various prompting strategies and four backbone models (closed-source models: GPT-3.5 and GPT-4; open-source models: Llama3 and QWen2.5) to identify the optimal ICL paradigm for CTI knowledge extraction, providing valuable insights for future research. CTINexus’s code and data are available at https://ctinexus.github.io/.
@@ 2.1 Cyber Threat Intelligence
Although crowd-sourced CTI reports provide valuable information, their unstructured format significantly hinders their effectiveness. As the number and complexity of cyberattacks increase, the textual CTI descriptions have also expanded, creating an urgent need for automated information extraction from CTI [72]. The extracted knowledge can be used to construct cybersecurity knowledge graphs (CSKGs), where nodes represent entities and edges represent relations. Compared to unstructured CTI text, CSKGs provide a holistic profile for cyber threats, offer better visualization, and are more amenable to integration into downstream applications. The construction of a CSKG typically follows an ontology, which specifies the entity types and their allowed relations. Despite the development of various security ontologies [49, 73, 79] covering different aspects of threats, the rapid evolution of threats makes it nearly impossible to maintain a universal, comprehensive ontology. This underscores the need for CTI knowledge extraction approaches that can adapt to different ontologies and emerging threats with minimal transition effort.
@@ 2.2 Limitations of Existing Approaches
  • Domain complexity: The grammatical rules can apply to any domain. However, CTI text has several peculiarities that can confuse syntax parsing, leading to inaccurate extraction. Cybersecurity entities can contain special characters, such as dots in IPv4 addresses, underscores in file names, and slashes in file paths. These special characters can confuse basic NLP modules, like sentence segmentation and tokenization, which syntax parsing relies upon.
@@
On the other hand, fine-tuning-based approaches fine-tune pre-trained neural networks on annotated CTI domain datasets to perform named entity recognition (NER) and relation extraction (RE). For example, EXTRACTOR [74] fine-tunes a pre-trained BERT [75] model with thousands of annotated CTI sentences, to perform semantic role labeling to extract subjects, objects, and verb actions. AttacKG [56] fine-tunes a pre-trained model in the SpaCy library [46] to recognize entities and extract dependencies. LADDER [32] fine-tunes different pre-trained transformers, including BERT, RoBERTa, and XML-RoBERTa, on their custom datasets annotated according to their own ontology for performing NER and RE. ThreatKG [42] trains domain-specific BiLSTM and PCNN-ATT models for extracting security entities and relations. However, fine-tuning-based methods also have several drawbacks:
  • Resource requirement: Model training and fine-tuning require large amounts of labeled data (i.e., annotated CTI text corpora), and the labeling needs to be aligned with the targeted ontology. Such annotations are expensive to obtain, especially for emerging threats. Additionally, fine-tuning can be computationally expensive if the backbone model contains lots of parameters.
  • Ontology lock-in: Since the models are fine-tuned on datasets annotated using a specific ontology, they are difficult to generalize to new ontologies that cover different entities and relations. Transferring to other ontologies would require reannotating vast data and retraining the models, which is very costly.
@@ 2.2.1 Motivating Example
As shown in Fig. 1, the CSKG constructed by CTINexus is comprehensive, well-connected, and has much better quality, addressing all previous drawbacks. By leveraging the in-context learning of LLMs, the construction of such a CSKG does not rely on large amounts of training data and can adapt to different ontologies. We describe our approach in Section 4.
Figure 2: Overview of CTINexus. CTINexus comprises three phases. Phase 1, Cybersecurity Triplet Extraction, enables end-to-end extraction of cybersecurity triplets using in-context learning of LLM. Phase 2, Hierarchical Entity Alignment, reduces the redundancy of CSKG through coarse-grained grouping and fine-grained clustering. Phase 3, Long-Distance Relation Prediction, connects disjoint subgraphs by identifying central nodes and performing relation inference.
2.3 Large Language Models (LLMs)
@@ 3 Overview
Fig. 2 illustrates CTINexus. CTINexus introduces an ICL-based approach for data-efficient CTI knowledge extraction and CSKG construction. Unlike previous methods, CTINexus eliminates the need for extensive data annotations and parameter tuning, facilitating generalization to various ontologies. CTINexus focuses on constructing connected and comprehensive CSKGs, enabling entity alignment and long-distance relation inference. CTINexus includes three phases.
@@ 4.2 Cybersecurity Triplet Extraction
Figure 3: Comparison of CTINexus’s ICL-based CTI knowledge extraction (left) and a multi-turn QA-based extraction (right). CTINexus consolidates task descriptions (including applied ontology), $k$ selected demonstrations, and query into a single instruction for efficient cybersecurity triplet extraction. In contrast, the multi-turn QA paradigm requires multiple rounds of conversations with multiple prompts to extract different entities and relations, which is inefficient.
Figure 4: The design of CTINexus’s hierarchical entity alignment. The coarse-grained entity grouping phase populates an ICL prompt to assign entity types to the extracted triplets according to the applied ontology. Entities with the same type are grouped together. The fine-grained entity merging phase then uses an embedding-based technique to merge semantically similar entities within each group based on a predefined similarity threshold. During this phase, IOC protection is enforced to prevent erroneously merging semantically similar but conceptually distinct IOC entities.
Given that CTI text may contain diverse relations and we want the approach to be adaptable to emerging threats, we formulate the cybersecurity triplet extraction module in our pipeline as a semi-open extraction problem: Entity types follow MALOnt, as its coverage is already comprehensive, while relation extraction is modeled as open RE to maximize the coverage. These approaches transform information extraction tasks into multi-turn question-answering, leveraging the conversational capabilities of LLMs. Fig. 3 illustrates this paradigm. This method involves creating multiple questioning prompts for each information type and refining the responses. However, applying this multi-turn QA formulation to cybersecurity entity and relation extraction requires numerous lengthy prompts due to the extensive cybersecurity ontology that could contain many entity classes. For $N$ entities in the input CTI, $\frac{N(N-1)}{2}$ prompts are needed to extract relations between identified entities, leading to repetitive content and significant token waste, hindering scalability. Additionally, the multi-turn paradigm suffers from confirmation bias [36], as LLMs may confirm with a non-existing relation after several rounds of dialogue. In Section 5.3, we present our evaluation of prompt formulation and strategy that underpin CTINexus’s superiority over the multi-turn QA formulation and baseline prompt designs.
@@
    Demonstrations: Top-$k$ most relevant examples are retrieved using the demonstration retriever. Each consists of a CTI report annotated with the security triplets. These examples are ordered in ascending similarity to the input query based on our findings described in Section 5.3.
@@ 4.3 Hierarchical Entity Alignment
Entity alignment identifies entities with different mentions that refer to the same real-world object, a key area in knowledge graph research [29]. Aligning these mentions integrates sub-graphs containing complementary knowledge, enhancing the comprehensiveness of the knowledge graph. Traditional entity alignment techniques rely on heuristics like string matching and structural similarities, which fail to capture the underlying semantics or context of entities and have limited accuracy. Recent studies [78, 81, 69] have adopted deep learning-based methods to learn vector representations (i.e., embeddings) of entities, achieving better accuracy. However, embedding-based techniques face unique challenges in our problem domain. In CTI text, entities with similar embeddings may refer to different concepts, e.g., “.akira files” (an IOC) and “Akira” (a threat actor). Besides, comparing the semantic distance between every pair of entities has a computational complexity of $n^{2}$, where $n$ is the total number of entities. This is inefficient when $n$ becomes large.
@@
To further ensure safe merging, we introduce an IOC protection mechanism that prevents merging semantically similar but conceptually distinct IOCs sharing the same general type (e.g., CVE-2023-23397 vs. CVE-2023-23392). Specifically, we use a curated set of regular expressions to identify IOC patterns (e.g., CVE strings, IP addresses, file hash, etc.) and isolate them from being merged with other mentions.